# The future of offshore wind power production: wake and climate impacts

### v1

@@ Abstract
Rapid deployment of offshore wind is expected within the coming decades to help meet climate goals. With offshore wind turbine lifetimes of 25–30 years, and new offshore leases spanning 60 years, it is vital to consider long-term changes in potential wind power resource at the farm planning stage. Such changes may arise from multiple sources, including climate change, and increasing wake-induced power losses. In this work, we investigate and compare these two sources of long-term change in wind power, for a case study consisting of 21 wind farms within the German Bight. Consistent with previous studies, we find a small but significant reduction in wind resource due to climate change by the end of the 21st century under the high-emission RCP8.5 scenario, compared with a historical period, with a mean power reduction (over an ensemble of seven climate models) of 2.1%. To assess the impact of wake-induced losses due to increasingly dense farm build-out, we model wakes within the German Bight region using an engineering wake model, under various stages of (planned) build-out corresponding to the years 2010–2027. By identifying clusters of wind farms, we decompose wake effects into long-range (inter-cluster), medium-range (intra-cluster) and short-range (intra-farm) effects. Inter-cluster wake-induced losses increase from 0 for the 2010 scenario to 2.5% for the 2027 scenario, with intra-cluster losses also increasing from 0 to 4.3%. Intra-farm losses are relatively constant, at around 13%. While the evolution of wake effects therefore outweighs the climate effect, and impacts over a shorter timescale, both factors are significant. We also find evidence of non-linear interactions between the climate and wake effects. Both climate change and evolving wake effects must therefore be considered within resource assessment and wind farm planning.
@@ 1.1 Wake effects
A number of wake modelling approaches exist, with different methods suited to different scales and applications (Amiri et al., 2024, Wang et al., 2024). CFD models (Sanderse et al., 2011) can provide high-fidelity simulations of wakes from individual turbines or within farms, but are computationally expensive and therefore not suited to applications requiring a large number of simulations, such as array layout optimisation, or to the large spatial domains required to model inter-farm effects. Numerical weather prediction (NWP) models are also capable of simulating wake effects, but the spatial resolution at which they are typically applied is best suited to inter-farm/cluster wake analysis (Pryor and Barthelmie, 2024b), while their computational cost is also too high for use within optimisation studies. So-called engineering wake models have been developed since the 1980s (Jensen, 1983, Katic et al., 1986), and remain a commonly-used approach for simulating the effect of wakes on turbine power output. Although engineering models may use a variety of methods including analytic approaches or lookup tables, they are characterised by their strong simplifying assumptions (e.g. spatially and temporally uniform flow) but low computational cost, and are thus commonly applied to layout optimisation and control problems (Azlan et al., 2021). On these scales, the simplifying assumptions hold fairly well and the low computational cost can be leveraged for efficient optimisation. The TurbOPark engineering wake model (Nygaard et al., 2020, Pedersen et al., 2022) has gained popularity, having been developed in an attempt to accurately capture wake decay over long (e.g. inter-farm/cluster) distances. The overestimation of wake recovery was a known issue with earlier engineering wake models (Stieren and Stevens, 2021, Fischereit et al., 2022), resulting in underestimation of inter-farm wake effects (Pedersen et al., 2022). TurbOPark has been previously validated using SCADA data (Nygaard et al., 2022) and LES simulations (Stieren and Stevens, 2021), showing good performance in capturing long-range interactions. It is for this reason that we employ the TurbOPark model within this study.
@@
Although some of the above studies include sensitivity analysis with respect to the types and densities of the turbines deployed, they still consist of only a single build-out stage, and do not focus on the time evolution of build-out. This time-evolution is central to long-term policy and decision making, and contributes to current regulatory gaps regarding the impact of inter-farm wakes (Lundquist et al., 2019, Finserås et al., 2024).
@@ 1.2 Climate change
While the impact of climate change on wind energy resources in the North Sea is reasonably well studied, including the need for bias correction, such studies typically neglect wind turbine and farm wakes. An exception is Hahmann et al. (2022), who investigated the impact of climate change on wind energy resource in the North Sea, utilising an ensemble of CMIP6 projections. The study included wind turbine wake effects based on a planned Energy Island cluster of wind farms, modelled using a simple engineering wake model (Jensen, 1983). They found a significant decrease in energy production during summer of 6.9% for the 2031–2050 period under the SSP5-8.5 scenario, of which 0.7% was due to wake effects.
@@ 1.3 Contributions of this study
Furthermore, the findings of Hahmann et al. (2022) described above suggest that there is potential benefit to further studies into the interaction between climate change and wind turbine wake effects.
@@
The remainder of the paper is structured as follows. The methods are presented in section 2, including a description of the study region, wind farm data, reanalysis and climate projection datasets, bias correction methods, and wake modelling approach, as well as a description of the experiments we perform. The corresponding results are presented and discussed in section 3. Finally, we draw conclusions in section 4.
@@ 2.1 Study region & wind farm data
Most wind turbine power curves are not publicly available. We instead generate plausible power curves using PyWake’s ‘generic_wind_turbines’ module (Pedersen et al., 2019), based on publicly available turbine data (Bauer and Matysik, 2011) (hub height, rotor diameter, cut-in speed, etc).
@@ 2.3 Climate projection data
In this study we utilise regional climate simulations from EURO-CORDEX (Jacob et al., 2014), accessed via Earth System Grid Federation (2024). These simulations are downscaled from CMIP5 climate models (Taylor et al., 2012), on a regular 0.11 resolution grid (or approximately 12 $\mathrm{km}$). Although CMIP5 has been superseded by CMIP6, the availability of downscaled CMIP6 projections remains limited, hence we use EURO-CORDEX projections. The specific EURO-CORDEX models we choose are summarised in table 2. These models were selected on the basis that they provide wind speed data at 100 $\mathrm{m}$ height, at 3-hourly temporal resolution. The same set of models have previously been evaluated for the purposes of wind power generation by Luzia et al. (2023).
The two time periods we consider are 1976–2006, and 2070–2100. The first period is derived from ‘historical’ EURO-CORDEX simulations, and for the latter we use data corresponding to the high-emission RCP8.5 scenario. This scenario is the most commonly studied in the literature (Jung and Schindler, 2022), but results should be interpreted in the context that it is expected to produce the most significant climate change signal. Our choice to use the RCP8.5 scenario is based on the availability of downscaled climate model outputs at our desired spatial and temporal resolutions.
@@ 2.6 Turbine power and wake modelling
The power and wake modelling within this study utilises the implementation of the TurbOPark model (Pedersen et al., 2022) available within the PyWake package (Pedersen et al., 2019). The wake expansion parameter is set to 0.06. This is different to its original recommended value of 0.04 (Pedersen et al., 2022), but produced better results when applied to ERA-5 data and compared with FINO1 observations (see section 2.8.1). This is consistent with van der Laan et al. (2023). We employ squared sum wake superposition. We neglect blockage effects in order to enable more computationally efficient simulation, but preliminary experiments including a blockage model showed negligible impact to the results. Since the climate models provide the wind speed at 100 $\mathrm{m}$, we assume a power law vertical profile with an exponent of 0.11, to extrapolate to turbine hub heights. A power law profile is commonly assumed for extrapolating wind speeds from 10 $\mathrm{m}$ to hub height (Tobin et al., 2015), and we expect errors to be smaller here due to extrapolating from 100 $\mathrm{m}$ data. The turbines within our case study have hub heights in the range 89.5–150 $\mathrm{m}$. The choice of an exponent of 0.11 is recommended in the literature for offshore applications (Hsu et al., 1994).
@@
For most results we present herein, we refer only to total wake-induced losses. These are calculated by running PyWake twice. Simulation PW1 includes no wake deficit model, and is equivalent to simply extrapolating wind speeds to hub height and passing them directly through the turbine power curves. Simulation PW4 includes the wake deficit model described above. The losses per turbine are defined as the difference in power estimated by simulations PW4 and PW1.
In some cases we attribute the total wake-induced losses to interactions at different scales. Specifically, we refer to the intra-farm, intra-cluster and inter-cluster wake-induced losses; see figure 2. These are estimated by running two additional sets of PyWake simulations. Simulation PW2 is a composite of 21 separate PyWake simulations, with one simulation per farm. The intra-farm wake-induced losses can be directly estimated from the difference between simulations PW2 and PW1. Simulation PW3 is a composite of eight separate PyWake simulations, with one simulation per cluster (see figure 1). This enables the estimation of intra-cluster effects (the difference between simulations PW3 and PW2), and inter-cluster effects (the difference between simulations PW4 and PW3).
@@ 3.3.2 Change in power output and wake effects
The unwaked power is more sensitive than the wake effects to climate change. For the all-season results, all models predict a decrease in mean power, of up to 4%. The spread of climate projections for individual seasons is much greater. All models predict decreasing power for spring and summer, but there is no consensus for autumn or winter. The increased significance of changes in summer is consistent with previous literature (Carvalho et al., 2017, Moemken et al., 2018). The overall predicted reduction in wind resource have potential consequences for (robust) future wind farm designs. For example, cost-effective wind farm design may require larger and lower density farms in order to compensate for reduced resource density.

### v2

@@ Abstract
Rapid deployment of offshore wind is expected within the coming decades to help meet climate goals. With offshore wind turbine lifetimes of 25–30 years, and new offshore leases spanning 60 years, it is vital to consider long-term changes in potential wind power resource at the farm planning stage. Such changes may arise from multiple sources, including climate change, and increasing wake-induced power losses. In this work, we investigate and compare these two sources of long-term change in wind power, for a case study consisting of 21 wind farms within the German Bight. Consistent with previous studies, we find a small but significant reduction in wind resource due to climate change by the end of the 21st century under the high-emission RCP8.5 scenario, compared with a historical period, with a mean power reduction (over an ensemble of seven climate models) of 2.1%. To assess the impact of wake-induced losses due to increasingly dense farm build-out, we model wakes within the German Bight region using an engineering wake model, under various stages of (planned) build-out corresponding to the years 2010–2027. By identifying clusters of wind farms, we decompose wake effects into long-range (inter-cluster), medium-range (intra-cluster) and short-range (intra-farm) effects. Inter-cluster wake-induced losses increase from 0 for the 2010 scenario to 2.5% for the 2027 scenario, with intra-cluster losses also increasing from 0 to 4.3%. Intra-farm losses are relatively constant, at around 13%. While the evolution of wake effects therefore outweighs the climate effect, and impacts over a shorter timescale, both factors are significant. We also find evidence of an interaction between the climate and wake effects. Both climate change and evolving wake effects must therefore be considered within resource assessment and wind farm planning.
@@ 1.1 Wake effects
A number of wake modelling approaches exist, with different methods suited to different scales and applications (Amiri et al., 2024, Wang et al., 2024). CFD models (Sanderse et al., 2011) can provide high-fidelity simulations of wakes from individual turbines or within farms, but are computationally expensive and therefore not suited to applications requiring a large number of simulations, such as array layout optimisation, or to the large spatial domains required to model inter-farm effects. Numerical weather prediction (NWP) models are also capable of simulating wake effects, but the spatial resolution at which they are typically applied is best suited to inter-farm/cluster wake analysis (Pryor and Barthelmie, 2024b), while their computational cost is also too high for use within optimisation studies. So-called engineering wake models have been developed since the 1980s (Jensen, 1983, Katic et al., 1986), and remain a commonly-used approach for simulating the effect of wakes on turbine power output. Although engineering models may use a variety of methods including analytic approaches or lookup tables, they are characterised by their strong simplifying assumptions (e.g. spatially and temporally uniform flow) but low computational cost, and are thus commonly applied to layout optimisation and control problems (Azlan et al., 2021). On these scales, the simplifying assumptions hold fairly well and the low computational cost can be leveraged for efficient optimisation.
[…]
@@
Although some of the above studies include sensitivity analysis with respect to the types and densities of the turbines deployed, they still consist of only a single build-out stage, and do not focus on the time evolution of build-out. This time-evolution is central to long-term policy and decision making, and contributes to current regulatory gaps regarding the impact of inter-farm wakes (Lundquist et al., 2019, Finserås et al., 2024). It is for these reasons that the present study includes an analysis of the time evolution of wake effects due to build-out, over a period of two decades. Furthermore, while existing studies typically focus on either intra-farm or inter-farm/cluster wake effects, here we distinguish between effects at different spatial scales, in order to gain insight into the development of wake losses with build-out.
@@ 1.2 Climate change
While the impact of climate change on wind energy resources in the North Sea is reasonably well studied, including the need for bias correction, such studies typically neglect wind turbine and farm wakes. An exception is Hahmann et al. (2022), who investigated the impact of climate change on wind energy resource in the North Sea, utilising an ensemble of CMIP6 projections. The study included wind turbine wake effects based on a planned Energy Island cluster of wind farms, modelled using a simple engineering wake model (Jensen, 1983). They found a significant decrease in energy production during summer of 6.9% for the 2031–2050 period under the SSP5-8.5 scenario, of which 0.7% was due to wake effects. These results suggest that the interaction between climate and wake effects warrants further investigation. Furthermore, any long-term studies of wind resource, spanning farm lifetimes or offshore lease lengths, may suffer from systematic errors unless both climate and wake effects are considered.
@@ 1.3 Contributions of this study
Long-term changes in wind resources due to climate change are, in contrast, very well studied. However, the findings of Hahmann et al. (2022) described above suggest that there is potential benefit to further studies into the interaction between climate change and wind turbine wake effects, and that omitting wake effects when assessing climate change may lead to misleading conclusions.
@@
The remainder of the paper is structured as follows. The methods are presented in section 2, including a description of the study region, wind farm data, reanalysis and climate projection datasets, bias correction methods, and wake modelling approach, as well as a description of the experiments we perform. The corresponding results are presented in section 3. A discussion of the results, including the limitations of this study and avenues for further work, can be found in section 4. Finally, we draw conclusions in section 5.
@@ 2.1 Study region & wind farm data
Most wind turbine power curves are not publicly available. We instead generate plausible power curves using PyWake’s ‘generic_wind_turbines’ module (Pedersen et al., 2019), based on publicly available turbine data (Bauer and Matysik, 2011) (nominal power, hub height, rotor diameter, cut-in and cut-out speeds). This requires further assumptions regarding mechanical and electrical losses, as well as the peak thrust coefficient. We find that the peak thrust coefficient is particularly influential (and uncertain), and we therefore include a sensitivity analysis with respect to this parameter in A.
@@ 2.3 Climate projection data
In this study we utilise regional climate simulations from EURO-CORDEX (Jacob et al., 2014), accessed via Earth System Grid Federation (2024). These simulations are downscaled from CMIP5 climate models (Taylor et al., 2012), on a regular 0.11 resolution grid (or approximately 12.5 $\mathrm{km}$). Although CMIP5 has been superseded by CMIP6, the availability of downscaled CMIP6 projections remains limited, hence we use EURO-CORDEX projections. The specific EURO-CORDEX models we choose are summarised in table 2. These models were selected on the basis that they provide wind speed data at 100 $\mathrm{m}$ height, at 3-hourly temporal resolution. The same set of models have previously been evaluated for the purposes of wind power generation by Luzia et al. (2023).
The time periods we consider are 1976–2006 and 2070–2100, and a further period of 2040–2070 is also analysed in B. The first period is derived from ‘historical’ EURO-CORDEX simulations, and for the latter two periods we use data corresponding to the high-emission RCP8.5 scenario. This scenario is the most commonly studied in the literature (Jung and Schindler, 2022), but results should be interpreted in the context that it is expected to produce the most significant climate change signal. Our choice to use the RCP8.5 scenario is based on the availability of downscaled climate model outputs at our desired spatial and temporal resolutions. In B we also present results from the RCP2.6 scenario, for the 2070–2100 period.
@@ 2.6 Turbine power and wake modelling
Wake effects are simulated within this study using the TurbOPark engineering wake model (Pedersen et al., 2022). This choice is motivated by the computational efficiency of such models, which is well suited to long-term studies, and essential to the analysis of wake effects at different length scales, which requires many simulations with different turbine configurations. The disadvantages of engineering wake models compared with alternative methods such as mesoscale numerical weather prediction (NWP) models, such as assumptions about the uniformity of the wind over the study region, are mitigated by the limited model domain extent within this study, making engineering wake models an appropriate choice. This is discussed further in section 4.
Specifically, this paper utilises the implementation of the TurbOPark model available within the PyWake package (Pedersen et al., 2019). The wake expansion parameter is set to 0.06. This is different to its original recommended value of 0.04 (Pedersen et al., 2022), but produced better results when applied to ERA-5 data and compared with FINO1 observations (see section 2.8.1). This is consistent with van der Laan et al. (2023). We employ squared sum wake superposition. We neglect blockage effects in order to enable more computationally efficient simulation, but preliminary experiments including a blockage model showed negligible impact to the results. Since the climate models provide the wind speed at 100 $\mathrm{m}$, we assume a power law vertical profile with an exponent of 0.11, to extrapolate to turbine hub heights. A power law profile is commonly assumed for extrapolating wind speeds from 10 $\mathrm{m}$ to hub height (Tobin et al., 2015), and we expect errors to be smaller here due to extrapolating from 100 $\mathrm{m}$ data. The turbines within our case study have hub heights in the range 89.5–150 $\mathrm{m}$. The choice of an exponent of 0.11 is recommended in the literature for offshore applications (Hsu et al., 1994).
@@
Figure 3: PyWake modelling workflow used for calculating wake losses at different spatial scales. A given simulation is based on decomposition of the model domain to a certain level. This work utilises simulations which decompose to turbine level (simulation PW1), farm level (PW2), cluster level (PW3) or no decomposition (PW4). The figure shows decomposition to the farm level. PyWake simulations use wind roses derived from downscaled climate projections from the EURO-CORDEX project, which have been bias-corrected using ERA-5 reanalysis data.
Throughout this work, we refer to wake losses at different spatial scales. Figure 3 outlines the modelling workflow we use to calculate these losses. To calculate total wake-induced losses for a given build-out stage (or set of farms), we perform this workflow twice. Simulation PW1 constitutes decomposition to turbine level, which is also equivalent to using no wake deficit model, or to simply extrapolating wind speeds to hub height and passing them directly through the turbine power curves. Simulation PW4 performs no decomposition, and constitutes a single PyWake simulation for the whole model domain, using the wake deficit model described above. The losses are defined as the difference in power estimated by simulations PW4 and PW1.
In some cases we attribute the total wake-induced losses to interactions at the intra-farm, intra-cluster and inter-cluster scales; see figure 2. For a given build-out stage, these are estimated by running two additional sets of PyWake simulations, each following the workflow of figure 3. For simulation PW2, we decompose to farm level, resulting in up to 21 separate PyWake simulations (one simulation per farm), with the total power then aggregated across these simulations. The intra-farm wake-induced losses can be directly estimated from the difference between simulations PW2 and PW1. Simulation PW3 is based on decomposition to cluster level, resulting in up to eight separate PyWake simulations (one simulation per cluster; see figure 1). This enables the estimation of intra-cluster effects (the difference between simulations PW3 and PW2), and inter-cluster effects (the difference between simulations PW4 and PW3).
@@ 3.3.2 Change in power output and wake effects
The unwaked power is more sensitive than the wake effects to climate change. For the all-season results, all models predict a decrease in mean power, of up to 4%. The spread of climate projections for individual seasons is much greater. All models predict decreasing power for spring and summer, but there is no consensus for autumn or winter. The increased significance of changes in summer is consistent with previous literature (Carvalho et al., 2017, Moemken et al., 2018). The overall predicted reduction in wind resource has potential consequences for (robust) future wind farm designs. For example, cost-effective wind farm design may require larger and lower density farms in order to compensate for reduced resource density.
@@ 3.4 Build-out effects
4 Discussion
The results of this study indicate that both climate change and wake effects constitute significant sources of long-term change in wind farm power output. Owing to farm lifetimes of 25–30 years, and offshore lease durations of 60 years, both of these factors should be considered at the wind farm planning stage.
The results of section 3.3.2 indicate that the magnitude of wake effects differs between the historical and future climates, implying an interaction between climate and wake effects. Looking across all seasons, all climate models agree that there is a reduction in mean wind speeds at this location, or equivalently an increase in frequency of lower speeds and a decrease in frequency of higher speeds. Wake effects are strongest within a fairly narrow wind speed range, which is dictated by the cut-in and rated speeds of the turbines. This means that changes in wake effects are determined by changes in the occurrence frequency of specific wind speeds. This leads to high sensitivity, and a lack of consensus between each climate model on the overall change in wake effects under the future climate scenario. However, isolating the summer months, all climate models agree that there will be an increase in the magnitude of wake-induced losses. This compounds the effects of seasonal variation (with lowest winds in summer), as well as the direct climate effect of weakening wind resources in the region. Given that wind farms are long-term investments, these results suggest that it is important to consider both climate change and wake effects at all spatial scales within wind farm planning.
The results of this study may inform the development of robust approaches to long-term farm planning and design. We note that there is limited agreement among the climate models in terms of which precise wind speeds and directions change in frequency in future. This lack of fine-scale agreement may prevent the design of specific farm layouts which are robust to future changes in wind climate, although it is likely that larger-scale design choices such as turbine sizes and densities can be informed by such projections, with lower densities and larger turbines likely to be favoured by reduced resource density in future. For questions of future wind farm planning and design, distinguishing between intra-farm, intra-cluster and inter-cluster wake effects is particularly useful. Intra-farm losses can always be mitigated via array layout optimisation when designing an individual farm (Piggott et al., 2022), and optimal control once the farm is operational. There may be some scope to reduce intra-cluster losses via similar mechanisms, particularly if multiple farms within the same cluster are being designed simultaneously and by the same developer, or will be operated cooperatively. From the perspective of a wind farm developer, inter-cluster losses will be the most challenging to mitigate, but from the perspective of policy-makers, these losses may be the most important for informing lease area valuation and long-term planning decisions.
We chose to use an engineering wake modelling approach within this work. The wind inputs to these models were wind roses derived from downscaled climate projections from the EURO-CORDEX project, with a resolution of 0.11, or around 12.5 $\mathrm{km}$. There are several advantages to this engineering wake modelling approach, which enable the analysis of key drivers of long-term change in wind power outputs. It is computationally efficient, which is essential to running the large number of simulations required to isolate wakes effects at different spatial scales, and the use of wind roses makes it straightforward to calculate average power and wake effects under different climate scenarios. A disadvantage of this method is the introduction of unknown hyperparameters, in this case the wake expansion parameter. Previous work has shown that the optimal value of this parameter may vary with wind conditions, and the spatial scale of the wakes being modelled (van Binsbergen et al., 2024). However, the comparisons we presented with observations from the FINO1 mast indicate that our model setup performs adequately. Alternative modelling approaches would include mesoscale NWP modelling, or CFD methods such as RANS or LES. Mesoscale modelling in particular is often used for long-term wind power assessments, and would have two main advantages over the engineering modelling approach. Firstly, the use of mesoscale models would further downscale the climate projection datasets, potentially producing a more accurate representation of the local wind conditions over the study site, and incorporating a wider variety of atmospheric variables into the estimation of wind farm power. Secondly, it would remove the assumption that wind fields are uniform over the model domain. Mesoscale spatial variations may be significant in study regions such as the German Bight (von Brandis et al., 2022). However, the computational cost of mesoscale NWP modelling would be prohibitive for the analysis performed within this work, where the focus is on long-term behaviour, and the wake analysis requires a large number of individual simulations with different turbine configurations. The use of engineering wake modelling within this study is therefore a practical choice, which is motivated by the extensive validation of the TurbOPark model in the literature (e.g.  Stieren and Stevens, 2021, Pedersen et al., 2022, Nygaard et al., 2022, Van Binsbergen et al., 2024, Sørensen et al., 2024, zum Berge et al., 2024), as well as the validation against FINO1 observations we performed within this study.
Assessing future changes due to either climate or wake effects carries significant uncertainty. Within this work, an ensemble of seven downscaled climate projections produces a range of responses in wind power for the study region, from around 0 to 4%. This is despite all of the climate projections relying on the same RCP8.5 emissions pathway. As described in B, simulations based on an alternative lower-emission scenario (RCP2.6) produce climate-induced changes in mean power which differ in magnitude compared with RCP8.5, although with a lack of agreement between the climate models as to which climate scenario produces the largest changes. Simulations for an intermediate future period (2040–2070) show a slight reduction in the mean climate impact compared with the 2070–2100 period, although again with a lack of agreement between different climate models. Overall, the results of B demonstrate the high uncertainty inherent in using climate projections to model future wind power production. With respect to wake effects, this study is based on known or planned turbine locations and capacities. Nevertheless, as demonstrated by the experiments summarised in A, uncertainty in precise turbine characteristics introduces a significant amount of uncertainty in the simulated wake losses. While this study has therefore demonstrated that changes in both climate and wake effects are impactful, further work is required to investigate the relative sizes of the uncertainties associated with each.
Uncertainty in wake effects will increase for build-out time horizons beyond those considered in this study. For example, future work could incorporate much larger wind farms likely to be developed in the coming decades (e.g. based on future scenarios such as Waldman et al. (2023)). However, the details of such scenarios, including their associated turbine design parameters, are highly speculative, and their simulation will require an appropriate analysis and treatment of the resulting uncertainties. Furthermore, as discussed above, expanding the study area may stretch the validity of the current modelling approach (which assumes spatially and temporally uniform wind fields), and instead require mesoscale modelling to correctly capture spatial variation across the domain. An intermediate option could be to use streamline-following engineering wake models (e.g. von Brandis et al., 2022). It is for these reasons that we chose to limit the present study to farms with known designs. Although this results in a discrepancy between the build-out and climate change timescales investigated, we are still able to draw conclusions about the relative impacts of these effects. Further exploration of the impact of future build-out at larger scales is left to future work.
For the study region of this paper, the change in wake-induced losses from two decades of build-out is of significantly greater magnitude than the change in power due to approximately 100 years of climate change under the RCP8.5 scenario. Therefore, while the impact of climate change on future wind power is well studied in the literature, the results of this study highlight the importance of long-term build-out as a key driver of change in wind power production. However, we also note that the German Bight is a region with a particularly high density of wind farms, while the impact of climate change on wind resources in the North Sea is relatively small compared with other regions worldwide (Warder and Piggott, 2024). There would be value in future work repeating the analysis of this study for other regions, where the relative magnitudes of the climate and wake effects will likely differ. However, this would rely on the availability of data specifying turbine locations, characteristics (power, rotor area, etc) and installation dates, which may be a limiting factor in extending this analysis globally.